OpenStudy (anonymous):

write each expression as the product of two binomials x^2+5x+6 x^2+5x-6 x^2+9x+20 2y^2+5y-3 Can someone please explain how to do this? Not just the answer :)

OpenStudy (ranga):

All it means is factor the expression. I will do the first, you try the rest. x^2+5x+6 = x^2 + 2x + 3x + 6 = x(x + 2) + 3(x + 2) = (x + 2)(x + 3)

OpenStudy (anonymous):

Thank you!

OpenStudy (ranga):

you are welcome.